# Padding Tone - Revealing the role of the padding tokens in T2I Models

### v1

@@ Abstract
Text-to-image (T2I) diffusion models rely on encoded prompts to guide the image generation process. Typically, these prompts are extended to a fixed length by adding padding tokens before text encoding. Despite being a default practice, the influence of padding tokens on the image generation process has not been investigated. In this work, we conduct the first in-depth analysis of the role padding tokens play in T2I models. We develop two causal techniques to analyze how information is encoded in the representation of tokens across different components of the T2I pipeline. Using these techniques, we investigate when and how padding tokens impact the image generation process. Our findings reveal three distinct scenarios: padding tokens may affect the model’s output during text encoding, during the diffusion process, or be effectively ignored. Moreover, we identify key relationships between these scenarios and the model’s architecture (cross or self-attention) and its training process (frozen or trained text encoder). These insights contribute to a deeper understanding of the mechanisms of padding tokens, potentially informing future model design and training practices in T2I systems.
@@
Michael Toker1     Ido Galil1,2     Hadas Orgad1     Rinon Gal2     Yoad Tewel2 Gal Chechik2,3     Yonatan Belinkov1 1Technion – Israel Institute of Technology        2NVIDIA        3Bar-Ilan University
@@ 1 Introduction
Text-to-image (T2I) models consist of two main components: a text encoder, which generates representations of the user’s prompt, and a diffusion model, which generates an image based on this representation. To standardize sequence lengths for efficient batch processing in training and inference, input prompts are padded to a fixed length with a special padding token. Unlike language models, where padding tokens are explicitly masked and thus ignored, the computation process of the T2I models can use these tokens as any other token. Despite their ubiquity, the potential impact of padding tokens on image generation outcomes has been overlooked.
We introduce two methods to evaluate the influence of tokens on different model components: (1) Intervention in the Text Encoder Output (ITE) and (2) Intervention in the Diffusion Process (IDP). Both methods build on causal mediation analysis, also known as activation patching (Imai et al., 2010; Vig et al., 2020; Zhang and Nanda, 2024). This technique involves perturbing specific inputs or intermediate representations to observe their effect on the output, helping to pinpoint the influential elements.
@@
In cases where our analysis with ITE indicates that padding tokens are not used by the text encoder, we further examine the role of padding tokens in the diffusion process. Particularly, we investigate whether significant information is written into the padding token representations throughout the diffusion process. Here we employ IDP, illustrated in Figure 8, to interpret the causal effect of the padding tokens during the diffusion process. We begin with a standard prompt padded to a fixed length, as well as an “only pads” prompt. However, in IDP, token replacement occurs before each attention block within the diffusion process and at every diffusion step. We repeat the procedure of selectively replacing either prompt tokens or padding tokens with clean pads, similarly to ITE. Figure 1 illustrates an example of images generated using this method.
We analyze six different T2I models and highlight two scenarios where padding tokens are utilized. First, when the text encoder was not frozen during training or fine-tuning, it learns to encode meaningful semantic information into these tokens. Second, in architectures with multi-modal attention mechanisms—such as Stable Diffusion 3 (Esser et al., 2024) and FLUX111blackforestlabs.ai—padding tokens carry meaningful information throughout the diffusion process, even if the text encoder itself does not directly encode it. Here, the padding tokens seem to act as “registers”, with information written into their representations to store and recall, similarly to findings from both language models and vision-language models (Darcet et al., 2024; Burtsev et al., 2020).
@@
    We find that T2I models with frozen text encoders ignore padding tokens. However, when the text encoder is trained or fine-tuned, padding tokens gain semantic significance.
@@
    We uncover that even when padding tokens are not utilized by the text encoder, for some architectures of the diffusion model, they can still function as “registers” and play a meaningful part in the diffusion process.
method
Figure 2: ITE: Interpreting information within pad tokens in the text encoder. We first encode the full prompt and an clean pads separately. Next, we keep the tokens we want to interpret and replace all other tokens with clean pad tokens. We then generate an image conditioned on this mixed representation. In the example shown here, we interpret the pad tokens in LLaMA-UNet, revealing semantic information embedded within the pad tokens.
@@ 2 Analysis of Padding in Text Encoding
In the T2I pipeline, the text encoder processes the input prompt $P=[P_{1},..,P_{k}]$, with a prompt length of $k$. To ensure a consistent input length, the prompt is usually padded to a fixed length, denoted as $N$. We denote this padded version of the prompt as $P_{\text{full}}$, which is a concatenation of the $k$ prompt tokens and the $N-k$ padding tokens:
@@
The text encoder then processes $P_{\text{full}}$, producing a constant-length encoded representation, which is subsequently used by the diffusion model for conditional image generation. We denote this encoded full prompt representation as $E_{\text{full}}$.
@@ 2.1 Method
Our goal is to evaluate the information encoded in the prompt-contextual padding tokens, and to measure their effect on the generated image. To do so, as illustrated in Figure 2, we generate images using partial representations of $E_{\text{full}}$ that isolate the effect of the padding tokens. We generate images based on modified representations of $E_{\text{full}}$ and compare them to images generated from the full prompt $E_{\text{full}}$. This enables us to visually express the information from different parts of the text input.
@@
These encoded pad tokens are then used in constructing the final mixed representation, which combines both the encoded prompt and clean padding tokens. We use the encoded pad tokens since they contain no information related to the current prompt, while maintaining the same length and distribution of the text encoder’s output. This allows us to effectively isolate the contribution of the padding tokens that are encoded alongside the full prompt tokens, helping us understand how much of the information in the final representation comes from the prompt itself versus the pads. Figure 3 demonstrated our method. First, we generate an image from the full prompt, which is how the image is generated in the standard pipeline (left column). Then, we generate an image that demonstrates the information in the non-pad tokens, by replacing the pad tokens with clean pads (middle column). Lastly, we generate an image demonstrating the information within the pad tokens, by replacing the non-pad tokens with clean pads (left column).
More formally, the mixed representation for generating an image from the prompt tokens only (middle column) is:
$E_{\text{prompt}}=\left[E_{\text{full}}^{0:k},E_{\text{clean}}^{k:N}\right],$ (2)
where $E_{x}^{i:j}$ represents the encoded tokens from index $i$ to $j$, and for a representation that generates an image from the prompt-contextual padding tokens only (right column):
$E_{\text{pads}}=\left[E_{\text{clean}}^{0:k},E_{\text{full}}^{k:N}\right]$ (3)
Figure 4: Average CLIP score over 5,000 images generated from the different representations: full prompt, only prompt, prompt-contextual pads and clean pads. LDM and LLaMA-UNet are the only models achieving high CLIP scores for images generated from padding tokens, indicating their use during text encoding. See Table 4 in the Appendix for standard deviations.
@@ 2.2 Experimental Setup
Our prompts are based on the Parti dataset (Yu et al., 2022), a benchmark containing over 1600 diverse and challenging prompts used to evaluate T2I models. To prevent using prompts that have leaked into the training corpus of the models, we select prompts from eight different challenge categories in Parti, and use GPT-4o222openai.com/index/hello-gpt-4o to generate an alternative set of prompts with similar style and complexity. We then manually review the prompts to ensure their coherence. This process results in 500 new prompts. The complete list of categories, along with the prompt used with GPT, can be found in Appendix A, and the full dataset is included in the supplementary material.
@@
To evaluate the generated images, we employ two key metrics: CLIP score (Hessel et al., 2021), which measures how well the generated images align with the prompts, and KID (Kernel Inception Distance) (Bińkowski et al., 2018), to evaluate the quality of generated images. KID is used to measure the similarity between the feature distributions of images generated from full representation and generated images after some causal intervention. Unlike FID (Heusel et al., 2017), which is based on Gaussian approximations, KID uses the maximum mean discrepancy (MMD) measure, making it more robust in practice, especially when dealing with smaller sample sizes.
@@ 2.3 Results
Figure 4 shows the average CLIP score over generations from different representations: “full”, “prompt”, “prompt-contextual pads” and “clean”. Stable Diffusion (versions 2+3) and FLUX models appear to make little to no use of padding tokens: CLIP scores for the full and prompt representations are nearly identical, while the prompt-contextual pads—containing only padding tokens—yield significantly lower scores. In contrast, LLaMA UNet and LDM contain significant semantic information in padding, with a higher CLIP score for the “prompt-contextual pads”, although the degradation in performance from “full” to “prompt” is small.
@@
Table 1: KID scores between the images generated from the prompt-contextual pads vs. images generated only from prompt representations. Lower is better. The KID is calculated w.r.t images generated from the full representation.
Text encoder training objective and its influence on padding usage
Our results suggest that the training objective of the text encoder significantly impacts how padding tokens are utilized. Many current T2I models, such as Stable Diffusion and FLUX, employ a frozen text encoder, with the diffusion model being trained on its encoded outputs. It may be that because the text encoder is not explicitly trained to process padding tokens for image generation, it does not effectively incorporate them during the textual encoding. As shown in Figure 4 and Table 1, in models that use frozen text encoders, images generated using the “prompt” representation yield the same CLIP score as those generated using the “full” representation, while images generated from “prompt-contextual padding” representations result in a very low CLIP score, almost as low as those generated from clean padding. In these models, the prompt KID is very high, meaning that the images are out of distribution. This suggests that in these models, the text encoder does not encode any meaningful information in the padding tokens, which makes them unnecessary for generating the final image.
Other models, like LDM and Lavi-Bridge, propose adapting the text encoder specifically for the image generation task. These methods train the text encoder, including the use of padding tokens, on the image generation objective, allowing it to effectively learn how to utilize padding. In these models, the results differ: images generated from full prompt tokens have lower scores compared to those generated using prompt representations, suggesting that the information encoded in the prompt tokens is insufficient to generate the correct images. Furthermore, images generated from the prompt-contextual padding tokens in these models yield much higher CLIP scores, even surpassing images generated from full prompt tokens in one of the models. KID of “prompt-contextual pads” in these models is comparatively low, indicating that the images generated from pads come from a closer distribution compared to the images generated from the full representation. Overall, this indicates that padding tokens play an important role in the text encoding process for image generation in these adapted models.
@@
Since LLaMA is a language model adapted for image generation using LoRa training, we can load the LoRa with a scaling factor, $\alpha$, to observe how gradually removing LoRa affects the number of used pad tokens. Our results in Figure 5 show that as $\alpha$ decreases, fewer pad tokens are used. This indicates that part of what the LoRa learns involves encoding information in more pad tokens.
@@ 3 Analysis of Padding in the Diffusion Process
Even when padding tokens contain no meaningful information after text encoding, the diffusion model might still make use of them during the diffusion process. To generate images from text prompts, T2I models use an attention mechanism to condition the generation process, typically following two common approaches: cross-attention and MM-DiT (Esser et al., 2024) blocks. In cross-attention, used in models like Stable Diffusion 2/XL, the model converts image patches into query vectors and text tokens into key and value vectors. The image patches gather information from the text based on an attention map, but the text representation remains unchanged throughout the process. In contrast, MM-DiT blocks, found in models like FLUX and Stable Diffusion 3, implement a multi-modal self-attention, by projecting both image patches and text token representations into query, key, and value vectors. Thus, both the image and text representations update and influence each other during the attention process. We therefore expect that models implementing cross-attention where the pads are not used in the text encoder would also not use them in the diffusion process. However, models implementing MM-DiT blocks can potentially aggregate information into the padding tokens, even if initially they contain no information.
@@
To explore this, we examine the attention maps between image patches and text representations, resulting in an attention map for each token (see example in Figure 7). While in Stable Diffusion XL only the prompt (and the end-of-text) tokens significantly attend to main areas in the image, in FLUX not only prompt tokens, but also many pad tokens contribute much attention to image areas (Figure 6). Moreover, generating images with FLUX and Stable Diffusion XL, with and without padding (Figure 11, App. E), reveals that FLUX without padding often misses key details, while Stable Diffusion XL remains consistent in its generations.

### v2

@@ Abstract
Text-to-image (T2I) diffusion models rely on encoded prompts to guide the image generation process. Typically, these prompts are extended to a fixed length by appending padding tokens to the input. Despite being a default practice, the influence of padding tokens on the image generation process has not been investigated. In this work, we conduct the first in-depth analysis of the role padding tokens play in T2I models. We develop two causal techniques to analyze how information is encoded in the representation of tokens across different components of the T2I pipeline. Using these techniques, we investigate when and how padding tokens impact the image generation process. Our findings reveal three distinct scenarios: padding tokens may affect the model’s output during text encoding, during the diffusion process, or be effectively ignored. Moreover, we identify key relationships between these scenarios and the model’s architecture (cross or self-attention) and its training process (frozen or trained text encoder). These insights contribute to a deeper understanding of the mechanisms of padding tokens, potentially informing future model design and training practices in T2I systems. [^1] Code available at padding-tone.github.io
@@
Michael Toker1     Ido Galil1,2     Hadas Orgad1     Rinon Gal2     Yoad Tewel2,4 Gal Chechik2,3     Yonatan Belinkov1 1Technion – Israel Institute of Technology        2NVIDIA        3Bar-Ilan University     4Tel-Aviv University {tok,orgad.hadas}@campus.technion.ac.il,  {idogalil.ig,yoad.tewel}@gmail.com belinkov@technion.ac.il
@@ 1 Introduction
Text-to-image (T2I) models consist of two main components: a text encoder, which generates representations of the user’s prompt, and a diffusion model, which generates an image based on this representation. To standardize sequence lengths for efficient batch processing in training and inference, input prompts are padded to a fixed length with a special padding token. Unlike language models, where padding tokens are explicitly masked and thus ignored, the computation process of the T2I models can use these tokens as any other token. Despite their ubiquity, the potential impact of padding tokens on image generation has been overlooked. Our goal in this work is to evaluate the influence of these tokens and determine whether the model learns to use semantically meaningless tokens.
We introduce two methods to evaluate the influence of tokens on different model components: (1) Intervention in the Text Encoder (ITE) and (2) Intervention in the Diffusion Process (IDP). Both methods build on causal mediation analysis, also known as activation patching (Imai et al., 2010; Vig et al., 2020; Zhang and Nanda, 2024). This technique involves perturbing specific inputs or intermediate representations to observe their effect on the output, helping to pinpoint the influential elements. Figure 1 illustrates one of our interventions, showing images generated with perturbations in different parts of the textual representation.
@@
In cases where our analysis with ITE indicates that padding tokens are not used by the text encoder, we further examine the role of padding tokens in the diffusion process. Particularly, we investigate whether significant information is written into the padding token representations throughout the diffusion process. Here we employ IDP, illustrated in Figure 9, to interpret the causal effect of the padding tokens during the diffusion process. We begin with a standard prompt padded to a fixed length, as well as an “only pads” prompt. However, in IDP, token replacement occurs before each attention block within the diffusion process and at every diffusion step. We repeat the procedure of selectively replacing either prompt tokens or padding tokens with clean pads, similarly to ITE.
We analyze six different T2I models and highlight two scenarios where padding tokens are utilized. First, when the text encoder was not frozen during training or fine-tuning, it learns to encode meaningful semantic information into these tokens. Second, in architectures with multi-modal attention mechanisms—such as Stable Diffusion 3 (Esser et al., 2024) and FLUX222blackforestlabs.ai—padding tokens carry meaningful information throughout the diffusion process, even if the text encoder itself does not directly encode it. Here, the padding tokens seem to act as “registers”, with information written into their representations to store and recall, similarly to findings from both language models and vision-language models (Darcet et al., 2024; Burtsev et al., 2020).
@@
    We find that T2I models with frozen text encoders (e.g., Stable Diffusion XL and Stable Diffusion 2) ignore padding tokens (Figure 2, first row). However, when the text encoder is trained or fine-tuned (LDM, LLama-UNet), padding tokens gain semantic significance (Figure 2, second row).
@@
    We uncover that even when padding tokens are not utilized by the text encoder, for some architectures with multi-modal self-attention in the diffusion model (Stable Diffusion 3 and FLUX), they can still function as “registers” and play a meaningful part in the diffusion process (Figure 2, last row).
Figure 2: The scenarios we observe: padding tokens may be effectively ignored (first row; image generated using ITE), affect the model’s output during text encoding (second row; image generated using ITE), or be used during the diffusion process (last row; image generated using IDP). Left: baseline. Right: our method.
Figure 3: ITE: Interpreting information within pad tokens in the text encoder. We first encode the full prompt and the clean pads separately. Next, we keep the tokens we want to interpret and replace all other tokens with clean pad tokens. We then generate an image conditioned on this mixed representation. In the example shown here, we interpret the pad tokens in LLaMA-UNet, revealing semantic information embedded within the pad tokens.
@@ 2 Analysis of Padding in Text Encoding
In the T2I pipeline, the text encoder processes the input prompt $P=[P_{1},..,P_{k}]$, $k$ tokens. To ensure a consistent input length, the prompt is usually padded to a fixed length, denoted as $N$. We denote this padded version of the prompt as $P_{\text{full}}$, which is a concatenation of the $k$ prompt tokens and the $N-k$ padding tokens:
@@
The text encoder then processes $P_{\text{full}}$, producing a constant-length encoded representation, which is subsequently used by the diffusion model for conditional image generation. We denote this encoded full prompt representation as $E_{\text{full}}$, which consists of the encoded prompt tokens and the encoded prompt-contextual padding tokens. [^3] One special case is the use of an EOS token in CLIP models, which is discussed in Appendix E.
@@ 2.1 Method
Our goal is to evaluate the information encoded in the prompt-contextual padding tokens, and to measure their effect on the generated image. To do so, as illustrated in Figure 3, we generate images using partial representations of $E_{\text{full}}$ that isolate the effect of the padding tokens. We generate images based on these partial representations of $E_{\text{full}}$ and compare them to images generated from the full prompt $E_{\text{full}}$. This enables us to visually express the information from different parts of the text input.
@@
These encoded padding tokens are then used in constructing the final mixed representation, which combines both the prompt-contextual tokens and clean padding tokens. We use the encoded padding tokens since they contain no information related to the current prompt, while maintaining the same length and distribution of the text encoder’s output. This allows us to effectively isolate the contribution of the padding tokens that are encoded alongside the full prompt tokens, helping us understand how much of the information in the final representation comes from the prompt itself versus the prompt-contextual padding tokens. Figure 4 demonstrated our method. First, we generate an image from the full prompt, which is how the image is generated in the standard pipeline (left column). Then, we generate an image that demonstrates the information in the non-pad tokens (e.g. encoded prompt tokens), by replacing the prompt-contextual padding tokens with clean pads (middle column). Lastly, we generate an image demonstrating the information within the prompt-contextual padding tokens, by replacing the non-pad tokens with clean pads (right column).
More formally, the mixed representation for generating an image from the encoded prompt tokens only (middle column) is:
$E_{\text{prompt}}=\left[E_{\text{full}}^{0:k},E_{\text{clean}}^{k+1:N-1}\right],$ (2)
where $E_{x}^{i:j}$ represents the encoded tokens from index $i$ to $j$. For a representation that generates an image from the prompt-contextual padding tokens only (right column):
$E_{\text{pads}}=\left[E_{\text{clean}}^{0:k},E_{\text{full}}^{k+1:N-1}\right]$ (3)
Figure 5: Average CLIP score over 5,000 images generated from the different representations: full prompt, only prompt, prompt-contextual pads and clean pads using ITE. LDM and LLaMA-UNet are the only models achieving high CLIP scores for images generated from padding tokens, indicating their use during text encoding. See Table 5 in the Appendix for standard deviations.
@@ 2.2 Experimental Setup
Our prompts are based on the Parti dataset (Yu et al., 2022), a benchmark containing over 1600 diverse and challenging prompts used to evaluate T2I models. To prevent using prompts that have leaked into the training corpus of the models, we select prompts from eight different challenge categories in Parti, and use GPT-4o444openai.com/index/hello-gpt-4o to generate an alternative set of prompts with similar style and complexity. We then manually review the prompts to ensure their coherence. This process results in 500 new prompts. The complete list of categories, along with the prompt used with GPT, can be found in Appendix A, and the full dataset is included in the supplementary material.
@@
To evaluate the generated images, we employ two key metrics: CLIP score (Hessel et al., 2021), which measures how well the generated images align with the textual prompt, and KID (Kernel Inception Distance) (Bińkowski et al., 2018), to evaluate the quality of generated images. KID is used to measure the similarity between the feature distributions of images generated from full representation and images generated after some causal intervention. Unlike FID (Heusel et al., 2017), which is based on Gaussian approximations, KID uses the maximum mean discrepancy (MMD) measure, making it more robust in practice, especially when dealing with smaller sample sizes.
@@ 2.3 Results
Figure 5 shows the average CLIP scores over generations from different representations: “full”, “prompt”, “prompt-contextual pads” and “clean”. Stable Diffusion (versions 2+3) and FLUX models appear to make little to no use of padding tokens: CLIP scores for the full and prompt representations are nearly identical, while the prompt-contextual pads—containing only padding tokens—yield significantly lower scores. In contrast, LLaMA UNet and LDM contain significant semantic information in padding, with a higher CLIP score for the “prompt-contextual pads”, although the degradation in performance from “full” to “prompt” is small.
@@
Table 1: KID scores between the images generated from the prompt-contextual pads vs. images generated only from prompt representations. All images are generated using ITE. Lower is better. The KID is calculated w.r.t. images generated from the full representation. Notably, LDM and LLaMA UNet are the only models that achieve a low KID on images generated from contextual pads (underlined).
Text encoder training objective and its influence on padding usage.
Our results suggest that the training objective of the text encoder significantly impacts how padding tokens are utilized. Many current T2I models, such as Stable Diffusion and FLUX, employ a frozen text encoder, with the diffusion model being trained on its encoded outputs. It may be that because the text encoder is not explicitly trained to process padding tokens for image generation, it does not effectively incorporate them during the textual encoding. As shown in Figure 5, in models that use frozen text encoders (Stable Diffusion and FLUX), images generated using the “prompt” representation yield the same CLIP score as those generated using the “full” representation, while images generated from “prompt-contextual padding” representations result in a very low CLIP score, almost as low as those generated from clean padding. Furthermore, Table 1 shows a clear distinction between models trained to process padding tokens for image generation and those that are frozen. For models with a frozen text encoder, the KID is very low (around 0.01) for images generated from prompt tokens only, whereas it is high for images generated from contextual padding tokens, indicating that these images are out of distribution. This suggests that in these models, the text encoder does not encode meaningful information in the padding tokens, making them unnecessary for generating the final image.
Other models, like LDM and Lavi-Bridge, adapt the text encoder specifically for the image generation task. These methods train the text encoder, including the use of padding tokens, on the image generation objective, allowing it to effectively learn how to utilize padding. In these models, the results differ: images generated from full prompt tokens have lower scores compared to those generated using prompt representations, suggesting that the information encoded in the prompt tokens is insufficient to generate the precise images. Furthermore, images generated from the prompt-contextual padding tokens in these models yield much higher CLIP scores, even surpassing images generated from prompt tokens in LLama-UNet. These models achieve relatively low KID on images generated from the prompt-contextual padding tokens, indicating that these images are close in distribution to those generated from the full representation. This suggests that in these models, the text encoder has learned to utilize the padding tokens during the textual encoding.
Overall, these results indicate that pads play an important role in the text encoding process for image generation in these adapted models.
@@
Since LLaMA is a language model adapted for image generation using LoRa training, we can load the LoRa with a scaling factor, $\alpha$, to observe how gradually removing LoRa affects the number of used pad tokens. Our results in Figure 6 show that as $\alpha$ decreases, fewer pad tokens are used. This indicates that part of what the LoRa learns involves encoding information into more pad tokens.
@@ 3 Analysis of Padding in the Diffusion Process
Padding tokens may not carry meaningful information after text encoding, yet some diffusion architectures may still utilize them during the diffusion process. To generate images from the encoded textual representation, T2I models use an attention mechanism to condition the generation process. This mechanism typically follows one of two common approaches: cross-attention and MM-DiT (Esser et al., 2024) blocks. In cross-attention, used in models like Stable Diffusion 2/XL, the model converts image patches into query vectors and text tokens into key and value vectors. The image patches gather information from the encoded textual representation, based on an attention map, but the text representation remains unchanged throughout the process. In contrast, MM-DiT blocks, found in models like FLUX and Stable Diffusion 3, implement a multi-modal self-attention, by projecting both image patches and text token representations into query, key, and value vectors. Thus, both the image and text representations update and influence each other during the attention process. We therefore expect that models implementing cross-attention where the pads are not used in the text encoder would also not use them in the diffusion process. However, models implementing MM-DiT blocks can potentially aggregate information into the padding tokens, even if initially they contain no information.
@@
To understand whether padding tokens are utilized during image generation, we examine the attention maps between image patches and text representations, resulting in an attention map between each text token and all the image tokens (see example in Figure 8). While in Stable Diffusion XL only the prompt (and the end-of-text) tokens significantly attend to main areas in the image, in FLUX not only prompt tokens, but also many pad tokens contribute much attention to main image areas (Figure 7). Moreover, generating images with FLUX and Stable Diffusion XL, with and without padding (Figure 12, Appendix G), reveals that FLUX without padding often misses key details, while Stable Diffusion XL remains consistent in its generations even without padding tokens.